# COLUMBUS: Evaluating COgnitive Lateral Understanding through Multiple-choice reBUSes

### v1

@@ 1 Introduction
Human problem-solving seamlessly combines vertical and lateral thinking (De Bono 2016). Vertical thinking is an analytical search process that rewards logic, rules, and rationality. It optimizes correctness by narrowing down on quality solutions and rejecting suboptimal ones (Hernandez and Varkey 2008). For example, resolving the question mark in Figure 1 (left) requires systematically identifying that all examples adhere to the formula: (left - top) $\times$ right + bottom = 77. Meanwhile, lateral thinking (De Bono 1971) is explorative, divergent, and creative (Hernandez and Varkey 2008). It expands the solution space by diverging into novel directions. As illustrated in the right part of Figure 1, visual, spatial, verbal, and numerical cues must be interpreted unconventionally (defying common sense (Jiang, Ilievski, and Ma 2024)), a process that lends itself to lateral thinking. In this example of a rebus puzzle, the numbers “$1$” phonetically represent the word “ONCE”, while the visual-spatial relationship between the blue letters “MO” and “ON” spell “BLUE MOON”. As “ONCE” is placed inside “BLUE MOON” this leads to the solution B) Once in a blue moon.
@@
    An experimental analysis with COLUMBUS of two human participants and representative state-of-the-art (SotA) vision-language models evaluated in a zero-shot setting, revealing that models perform decently but lag behind humans. Moreover, models benefit from human-curated descriptions but even the SotA ones struggle to generate such representations at the right level of abstraction.
@@ 3.2 Puzzle Rendering
Rebus puzzles include elements (i.e., text or icons) whose appearance and position are determined by latent rules triggered by specific keywords in the puzzle’s answer. [^2] See Appendix A.4 for the complete list of rules with keywords. This is illustrated in Figure 1 (right), where the words “ONCE”, “IN”, “BLUE”, and “MOON” determine the puzzle’s elements and their arrangement. We expect that SotA generative models cannot be reliably applied to generate rebus puzzles, a hypothesis that we validate in Section 6.5. Instead, we render a puzzle by a taxonomy-driven transformation of its input elements, which first produces a graph and subsequently an image (green part in Figure 2).
@@
Image Generation. We select one of four predefined templates using only the graph as input. The template selection is based on heuristics over the relational rules, as these define the positional properties of each element.
@@ 4 The COLUMBUS Benchmark
Quality Control. The graph parsing for all phrases and compounds includes a preprocessing step to remove stopwords that do not belong to the set of rule-triggering keywords. [^4] E.g., “to” is a stopword, but triggers the repetition: two rule phonetically. Appendix A.3 has the complete list of stopwords. Multiple elements with individual rules can still be present in the same puzzle, and more than one modifier rule can be applied to an element. However, we apply at most one individual rule to a single element. In cases where multiple individual rule can be applied to a single element, we generate these individually for each rule as separate puzzles. To further improve readability and limit the risk of overlapping elements, we restrict the image’s rendered elements using heuristics based on the number of elements and their rules (see Appendix A.5). We take the top 1,000 puzzles from the remaining puzzles with the most edges and rules per node to ensure the benchmark is challenging (see more details in Appendix A.1). To provide a fairer comparison between textual and icon puzzles, each puzzle containing an icon is duplicated, and all its icons are converted to their textual counterparts. Finally, we filter out low-quality puzzles with overlapping or overflowing text from this remaining set.
Benchmark Composition. We split the benchmark into two partitions: COLUMBUS-text that only contain text and COLUMBUS-icon that contain at least one icon. Between these two partitions, COLUMBUS features an overlap subset of 339 puzzle pairs. Each pair consists of two versions of the same puzzle: one version uses icons, and the other uses text instead of those icons. Table 1 shows key statistics about COLUMBUS. While non-icon puzzles are more numerous, icon puzzles feature more elements. This can be attributed to the difference in the answer length, as longer answers feature more chances that a word can be replaced with an icon. Specific rule categories, such as relational: next to and modifier: sound, are very prominent in the data, whereas others, such as size: big and relational: outside, have fewer data points. For the rule distributions in COLUMBUS, see Table 5 in Appendix A.6.
@@
General Number of puzzles 638 370 1,008
Mean answer length (# words) 4.12 5.35 4.57
Rules Freq. of individual rules 255 76 331
Freq. of relational rules 541 424 965
Freq. of modifier rules 505 255 760
Graphs Freq. of single node graphs 200 35 235
Freq. of double node graphs 333 245 578
@@
Distractors % of questions with distractors containing visible puzzle words 89.03 97.57 92.16
@@ 5 Experimental Design
As instruction-tuned models, we include 1) BLIP-2 coupled with Flan-T5-11b (Chung et al. 2022), which achieves SotA performance on zero-shot VQA tasks; 2) InstructBLIP (Dai et al. 2023), an instruction-tuned version of BLIP-2 model that uses Vicuna-7b (Zheng et al. 2023); 3) QwenVL (Bai et al. 2023b), a 7 billion visual multimodal version of the Qwen LLM (Bai et al. 2023a) from which we use the chat variant; 4) CogVLM (Wang et al. 2023), a 17 billion parameter VLM that achieves SotA performance on several VQA benchmarks; 5) Llava (Liu et al. 2023a), a large VLM that achieves SotA performance on several vision benchmarks despite its lack of billion-scale data. For Llava, we use the 13b (v1.5) and 34b (v1.6) variants, in which the 34b model is quantized to 4-bit; 6) Mistral-7b (v2) (Jiang et al. 2023a) to use in text-only, question-answering (QA) auxiliary experiments.
@@
For closed-source models, we selected four models from two representative families based on their promising performance in public visual reasoning benchmarks (Lu et al. 2023; Liu et al. 2023b): 1) GPT-4o and GPT-4o-mini (OpenAI et al. 2024) and 2) Gemini 1.5 (Pro) and Gemini 1.5 (Flash) (Anil et al. 2023).
@@
We experiment with two structural variants of closed-source models: forward and backward chaining (Jurafsky and Martin 2009). In forward chaining (FC), the model constructs evidence from an image and connects this evidence with the optimal candidate answer (Wang et al. 2024). The forward chaining approach first prompts models to generate JSON files with attributes (e.g., name, relation, description) for each visible puzzle element, which can later be used as a reference in the final prompting. As a representative of backward chaining (BC) from question and image towards the answer, we test belief graphs (Kassner et al. 2023) where a model recursively derives and evaluates explanations for each candidate answer. Belief graphs excavate additional information by recursively evaluating the truth assignments of premises generated for each answer candidate. The truth assignments are then optimized with a SAT solver, yielding the most probable answer. This approach is evaluated on a random subset of 50 puzzles (see Appendix B).
Human Evaluation.
@@
To estimate human performance on COLUMBUS, we ask two participants to answer a subset of 105 randomly selected puzzles (39 text and 40 icon puzzles, and 13 overlap puzzles with both a textual and icon variant). For more details about the setup, see Appendix D.
Model Inputs.
@@
We explore four human-curated input levels, each providing the model with increasing information about the puzzle, i.e., its description and details on the nodes or edges of a puzzle’s graph. Specifically: 1. no description of the nature of the puzzle, nor the graph; 2. only a description of the nature of the puzzle; 3. description of the nature of the puzzle and the graph nodes; 4. description of the nature of the puzzle and the full graph (both nodes and edges). See Table 6 in Appendix C.1 for the exact prompts used.
Evaluation Protocol.
@@
Following other multiple-choice benchmarks (Jiang et al. 2023b; Zhu et al. 2016; de Faria et al. 2023), we use accuracy as the evaluation metric, defined as the percentage of puzzles solved correctly. To extract answers from a model’s output, we use regex to check for choice symbols (e.g., “A”) if they are present and then perform exact string matching to the correct answer/symbol. For the larger, more flexible models that produce long explanations for their answers, we use GPT-4o to extract their answers automatically. For model outputs that answer a given puzzle with multiple options, we pick one of them randomly. We also compute the most frequent option for each model to check for imbalanced answer distributions.
@@ 6 Results
Table 2: Results for each model on COLUMBUS-text and COLUMBUS-icon. The accuracy and the most frequent answer (MFA) are reported based on a single run. Highest and second highest model results are highlighted in bold and underlined, respectively. The prompt includes a description of the nature of the puzzle (i.e., prompt 2). We did not test backward chaining for the Gemini models, as they do not output probabilities.
@@
Acc. MFA, % Acc. MFA, %
CLIP 56.3 D, 27.3 52.7 A, 26.5
BLIP-2 OPT (2.7b) 21.8 A, 44.3 24.3 A, 51.6
BLIP-2 OPT (6.7b) 24.1 A, 96.5 25.7 A, 95.1
Fuyu (8b) 32.0 C, 46.9 31.1 C, 43.2
InstructBLIP Vicuna (7b) 51.3 D, 58.8 51.4 D, 54.9
Qwen-VL (7b) 57.8 B, 29.0 64.3 B, 33.0
BLIP-2 Flan-T5-XXL (11b) 68.2 A, 34.9 72.2 A, 35.4
Llava (13b) 58.0 A, 31.1 58.9 A, 31.6
CogVLM (17b) 59.3 B, 35.9 60.3 B, 42.7
Llava (34b) 52.0 A, 29.9 49.5 B, 32.3
GPT-4o 81.4 A, 26.8 84.9 B, 29.1
GPT-4o-mini 74.8 B, 29.4 77.4 B, 31.3
Gemini 1.5 (Pro) 66.5 D, 31.3 70.4 D, 34.8
Gemini 1.5 (Flash) 61.6 C, 32.5 63.4 C, 34.4
GPT-4o (FC) 82.9 D, 28.8 79.8 D, 30.1
GPT-4o-mini (FC) 74.3 C, 27.7 72.8 D, 27.4
Gemini 1.5 (Pro) (FC) 74.2 D, 34.6 76.3 D, 33.0
Gemini 1.5 (Flash) (FC) 70.0 D, 29.5 80.9 C, 30.1
GPT-4o (BC) 61.3 A, 25.8 68.4 B, 42.1
GPT-4o-mini (BC) 38.7 A, 32.3 63.2 B, 42.1
Human 99.0 N/A 91.5 N/A
@@ 6.1 Overall Performance
Table 2 shows the performance of each model on COLUMBUS-text and COLUMBUS-icon. The closed-source and larger open-source models perform best on both partitions, while the small, non-instruction-tuned models perform near-randomly and have strong biases to certain answers (typically A or D). Comparing the average model performance with text and icons, we see no significant difference. Namely, the performance is slightly higher on COLUMBUS-icon, whereas on the overlapping set of 339 puzzles, we observe a slightly higher performance on the textual puzzles. The best model for each partition is consistently GPT-4o, which is the expected result. Yet, none of the models surpass human accuracy, with average gaps of 45.15% on COLUMBUS-text and 35.75% on COLUMBUS-icon.
Figure 5: Results for four prompts that supply the model with increasing information for COLUMBUS-text (left) and COLUMBUS-icon (right). The best-performing model from the following types is shown: open-source non-instruction VLM (Fuyu-8b), open-source instruction VLM (BLIP-2 Flan-T5-XXL), closed-source VLM (GPT-4o), and text-only LLM (Mistral-7b), as well as human accuracy.
@@ 6.2 Impact of Structural Reasoning
The two structured variants show opposing results (Table 2). Forward chaining, leading the model to generate graph descriptions in JSON format, yields marginal improvement on GPT4 while boosting Gemini performance greatly (6% to 17%). The structured JSON data allows Gemini to reason about rebuses with fine-grained detail, enabling the model to focus first on visual components and then handle broader reasoning. Both models achieve around 80% with forward chaining, leaving a gap against human performance. Here, we note that the forward chaining models benefit from including rule-related details in the metadata (e.g., a ’reverse’ label for words written backward), which motivates further work toward steering models to generate puzzle-specific abstractions. On the contrary, backward chaining yields an 18.3% and 25.15% drop in accuracy for GPT-4o and GPT-4o-mini, averaged across the two partitions. We ascribe this to the models lacking a global overview of the image, as each evaluated premise focuses on local parts of a puzzle without cohesively pointing towards a candidate answer.
@@ 6.3 Model Sensitivity to Input Information
Can models benefit from a ground-truth structured description of the puzzle provided in their input? Figure 5 shows that adding information about the nature of the puzzle (prompt 2) has little effect (+2.56% and -0.44% for textual and icon puzzles, respectively). Adding a description of the graph nodes (prompt 3) increases the model performance by $12.25\%$ and $14.8\%$ for non-icon and icon puzzles, respectively, reaching over 90% for GPT-4o. However, adding information on the relational rules only increases performance $4.12\%$ and $2.73\%$ for COLUMBUS-text and -icon, respectively. Interestingly, the forward chaining variant, inspired by prompt 3, also benefits from including information; however, the improvement is much higher when using ground-truth data, indicating that the models struggle to generate a graph description at the right level of abstraction. Considering the example in Figure 1, the models extract the text as is (e.g., extract “MO1111ON”) and cannot make the lateral connection that words/letters need rearranging. Even GPT-4o struggles consistently with this, such as with certain direction rules (see Section 6.4). For the results from the other models, see Table 8 in Appendix E.
Figure 6: Percentage of puzzles solved by GPT-4o in COLUMBUS-text and COLUMBUS-icon for each rule. It uses a prompt that only describes the nature of the puzzle (prompt 2). For COLUMBUS-icon, the direction rules are omitted because these either do not function with icons or become functionally identical to other rules when combined with icons.

### v2

@@ 1 Introduction
Human problem-solving seamlessly combines vertical and lateral thinking (De Bono 2016). Vertical thinking is an analytical search process that rewards logic, rules, and rationality. It optimizes correctness by narrowing down on quality solutions and rejecting suboptimal ones (Hernandez and Varkey 2008). For example, resolving the question mark in Figure 1 (left) requires systematically identifying that all examples adhere to the formula: (left - top) $\times$ right + bottom = 77. Meanwhile, lateral thinking (De Bono 1971) is explorative, divergent, and creative (Hernandez and Varkey 2008). It expands the solution space by diverging into novel directions. As illustrated in the right part of Figure 1, visual, spatial, verbal, and numerical cues must be interpreted unconventionally (defying common sense; Jiang, Ilievski, and Ma 2024), a process that lends itself to lateral thinking. In this example of a rebus puzzle, the numbers “$1$” phonetically represent the word “ONCE”, while the visual-spatial relationship between the blue letters “MO” and “ON” spell “BLUE MOON”. As “ONCE” is placed inside “BLUE MOON” this leads to the solution B) Once in a blue moon.
@@
    An experimental analysis with COLUMBUS with human participants and representative state-of-the-art (SotA) vision-language models evaluated in a zero-shot setting, revealing that models perform decently but lag behind humans. Moreover, models benefit from human-curated descriptions, but even the SotA ones struggle to generate representations at the right level of abstraction.
@@ 3.2 Puzzle Rendering
Rebus puzzles include elements (i.e., text or icons) whose appearance and position are determined by latent rules triggered by specific keywords in the puzzle’s answer. This is illustrated in Figure 1 (right), where the words “ONCE”, “IN”, “BLUE”, and “MOON” determine the puzzle’s elements and their arrangement. We expect that SotA generative models cannot be reliably applied to generate rebus puzzles, a hypothesis we validate in Section 6.5. Instead, we render a puzzle by a taxonomy-driven transformation of its input elements, which first produces a graph and subsequently an image (green part in Figure 2).
@@
Image Generation. We select one of four templates using only the graph as input. Templates include $x,y$ coordinates and a font-size multiplier. Three templates position up to three points equally along the $x$-axis in the image center, mapping graph nodes (left to right) to template points. The fourth handles graphs with the above rule. We change the appearance of the elements in an image according to the attributes of that element’s respective node.
@@ 4 The COLUMBUS Benchmark
Quality Control. The graph parsing for all phrases and compounds includes a preprocessing step to remove stopwords that do not belong to the set of rule-triggering keywords. [^3] E.g., “to” is a stopword, but triggers the repetition: two rule phonetically. Multiple elements with individual rules can still be present in the same puzzle, and more than one modifier rule can be applied to an element. However, we apply at most one individual rule to a single element. In cases where multiple individual rules can be applied to a single element, we generate these individually for each rule as separate puzzles. To further improve readability and limit the risk of overlapping elements, we restrict the image’s rendered elements using heuristics based on the number of elements and their rules. These exclude images generated on graphs that (1) have more than three nodes connected by the next to rule, (2) have an above rule where the top/bottom exceeds two nodes, (3) have an above rule where either the top/bottom exceeds two nodes, (4) have an outside rule where either the inside/outside portion has more than one node. We take the top 1,000 from the remaining puzzles with the most edges and rules per node to ensure the benchmark is challenging. To provide a fairer comparison between textual and icon puzzles, each puzzle containing an icon is duplicated, and all its icons are converted to their textual counterparts. Finally, we filter out low-quality puzzles with overlapping or overflowing text from this remaining set.
Benchmark Composition. We split the benchmark into two partitions: COLUMBUS-text that only contain text and COLUMBUS-icon that contain at least one icon. Between these two partitions, COLUMBUS features an overlap subset of 338 puzzle pairs. Each pair consists of two versions of the same puzzle: one version uses icons, and the other uses text instead of those icons. Table 1 shows key statistics about COLUMBUS. While non-icon puzzles are more numerous, icon puzzles feature more elements. This can be attributed to the difference in the answer length, as longer answers feature more chances that a word can be replaced with an icon.
@@
General Number of puzzles 634 371 1,005
Mean answer length (# words) 4.12 5.35 4.58
Rules Freq. of individual rules 253 76 329
Freq. of relational rules 540 425 965
Freq. of modifier rules 503 255 758
Graphs Freq. of single node graphs 197 35 232
Freq. of double node graphs 332 246 578
@@
Distractors % of questions with distractors containing visible puzzle words 89.27 97.57 92.34
@@ 5 Experimental Design
As instruction-tuned models, we include 1) BLIP-2 coupled with Flan-T5-11b (Chung et al. 2022), which achieves SotA performance on zero-shot VQA tasks; 2) InstructBLIP (Dai et al. 2023), an instruction tuned version of BLIP-2 model that uses Vicuna-7b (Zheng et al. 2023); 3) QwenVL (Bai et al. 2023b), a 7 billion visual multimodal version of the Qwen LLM (Bai et al. 2023a) from which we use the chat variant; 4) CogVLM (Wang et al. 2023), a 17 billion parameter VLM that achieves SotA performance on several VQA benchmarks; 5) Llava (Liu et al. 2023a), a large VLM that achieves SotA performance on several vision benchmarks despite its lack of billion-scale data. For Llava, we use the 13b (v1.5) and 34b (v1.6) variants; 6) Mistral-7b (v2) (Jiang et al. 2023a) to use in text-only, question-answering (QA) auxiliary experiments.
@@
For closed-source models, we selected four models from two representative families based on their promising performance in public visual reasoning benchmarks (Lu et al. 2023; Liu et al. 2023b): 1) GPT-4o and GPT-4o-mini (OpenAI 2024) and 2) Gemini 1.5 (Pro) and Gemini 1.5 (Flash) (DeepMind 2023).
@@
We experiment with two structural variants of closed-source models: forward and backward chaining (Jurafsky and Martin 2009). In forward chaining (FC), the model constructs evidence from an image and connects this evidence with the optimal candidate answer (Wang et al. 2024). The forward chaining approach first prompts models to generate JSON files with attributes (e.g., name, relation, description) for each visible puzzle element, which can later be used as a reference in the final prompting. As a representative of backward chaining (BC) from question and image towards the answer, we test belief graphs (Kassner et al. 2023) where a model derives and evaluates explanations for each candidate answer. Belief graphs excavate additional information by recursively evaluating the truth assignments of premises generated for each answer candidate. The truth assignments are then optimized with an SAT solver, yielding the most probable answer. This approach is evaluated on a random subset of 50 puzzles.
@@
Human Evaluation. To estimate human performance on COLUMBUS, we ask five participants to answer a subset of 103 randomly selected puzzles, consisting of 37 text, 40 icon puzzles, and 13 overlap puzzles with both a textual and icon variant.
@@
Model Inputs. We explore four human-curated input levels, each providing the model with increasing information about the puzzle, i.e., its description and details on the nodes or edges of a puzzle’s graph. Specifically: 1. no description of the nature of the puzzle, nor the graph; 2. only a description of the nature of the puzzle; 3. description of the nature of the puzzle and the graph nodes; 4. description of the nature of the puzzle and the full graph (both nodes and edges).
@@
Evaluation Protocol. Following other multiple-choice benchmarks (Jiang et al. 2023b; Zhu et al. 2016; de Faria et al. 2023), we use accuracy as the evaluation metric, defined as the percentage of puzzles solved correctly. To extract answers from a model’s output, we use regex to check for choice symbols (e.g., “A”) if they are present and then perform exact string matching to the correct answer/symbol. For the larger, more flexible models that produce long explanations for their answers, we use GPT-4o to extract their answers automatically. For model outputs that answer a given puzzle with multiple options, we pick one of them randomly. All models are run three times, and their performance is averaged to account for randomness.
@@ 6 Results
Table 2: Results for each model on COLUMBUS-text and COLUMBUS-icon. The accuracy’s mean and standard deviation (SD) are reported across three runs. The highest and second highest model results are highlighted in bold and underlined, respectively. The prompt includes a description of the nature of the puzzle (i.e., prompt 2). We did not test backward chaining for the Gemini models, as they do not output probabilities.
@@
Mean SD Mean SD
CLIP 56.15 0.00 52.56 0.00
BLIP-2 OPT (2.7b) 24.74 0.21 24.08 0.13
BLIP-2 OPT (6.7b) 23.95 0.16 25.61 0.00
Fuyu (8b) 32.02 0.00 31.00 0.00
InstructBLIP Vicuna (7b) 51.47 0.13 51.75 0.38
Qwen-VL (7b) 58.02 0.35 63.16 0.55
BLIP-2 Flan-T5-XXL (11b) 68.24 0.07 71.97 0.00
Llava (13b) 58.02 0.09 58.76 0.00
CogVLM (17b) 59.28 0.09 60.11 0.00
Llava (34b) 66.82 0.73 73.13 1.41
GPT-4o 80.89 0.97 83.34 1.11
GPT-4o-mini 73.96 0.71 77.69 0.49
Gemini 1.5 (Pro) 71.56 3.71 77.52 5.08
Gemini 1.5 (Flash) 64.42 2.00 67.44 2.93
GPT-4o (FC) 81.28 1.15 79.20 0.78
GPT-4o-mini (FC) 73.53 0.79 74.36 1.29
Gemini 1.5 (Pro) (FC) 69.98 3.00 72.10 3.17
Gemini 1.5 (Flash) (FC) 72.00 1.42 75.88 3.55
GPT-4o (BC) 64.37 1.63 71.67 4.71
GPT-4o-mini (BC) 45.93 8.65 60.00 4.08
Human 98.00 N/A 93.21 N/A
@@ 6.1 Overall Performance
Table 2 shows the performance of each model on COLUMBUS-text and COLUMBUS-icon. The closed-source and larger open-source models perform best on both partitions, while the small, non-instruction-tuned models perform near-randomly. Comparing the mean model performance with text and icons, we see no significant difference. Namely, the accuracy is slightly higher on COLUMBUS-icon, whereas on the overlapping set of 338 puzzles, we observe a slightly higher accuracy on the textual puzzles. As expected, the best model for each partition is consistently GPT-4o. Yet, none of the models surpass human accuracy, with average gaps of 38.17% on COLUMBUS-text and 30.64% on COLUMBUS-icon.
Figure 5: Results for four prompts that supply the model with increasing information for COLUMBUS-text (left) and COLUMBUS-icon (right) (averaged across three runs). The best-performing model from the following types is shown: open-source non-instruction VLM (Fuyu-8b), open-source instruction VLM (BLIP-2 Flan-T5-XXL), closed-source VLM (GPT-4o), and text-only LLM (Mistral-7b), as well as human accuracy.
@@ 6.2 Impact of Structural Reasoning
The two structured variants show opposing results (Table 2). Forward chaining, leading the model to generate graph descriptions in JSON format, yields little effect on the performance of GPT-4o (-1.88%) and Gemini (+2.26%), averaged across both models and partitions. Both models suffer from a gap against human performance. On the contrary, backward chaining yields an 14.1% and 22.86% drop in accuracy for GPT-4o and GPT-4o-mini averaged across the two partitions. We ascribe this to the models lacking a global overview of the image, as each evaluated premise focuses on local parts of a puzzle without cohesively pointing towards a candidate answer.
@@ 6.3 Model Sensitivity to Input Information
Can models benefit from a ground-truth structured description of the puzzle provided in their input? Figure 5 shows that adding information about the nature of the puzzle (prompt 2) has little effect (+2.68% and +3.39% for textual and icon puzzles, respectively). Adding a description of the graph nodes (prompt 3) increases the model performance by $11.91\%$ and $14.9\%$ for non-icon and icon puzzles, respectively, reaching over 90% for GPT-4o. However, adding information on the relational rules only increases performance $1.47\%$ and $0.56\%$ for COLUMBUS-text and -icon, respectively. Considering the example in Figure 1, the models extract the text as is (e.g., extract “MO [^1] ON”) and cannot make the lateral connection that words/letters need rearranging. Even GPT-4o struggles consistently with this, such as with certain direction rules, as discussed in the next Section.
Figure 6: Percentage of puzzles solved by GPT-4o for a single run in COLUMBUS-text and COLUMBUS-icon for each rule. The prompt used only describes the nature of the puzzle (prompt 2). For COLUMBUS-icon, the direction rules are omitted because these either do not function with icons or become functionally identical to other rules when combined with icons.